OpenStudy (anonymous):

find the exact value of sin(cos inverse(-3/7)+ tan inverse (9/4))

OpenStudy (anonymous):

\[\sin(\cos^{-1} (-3/7)+ \tan^{-1} (9/4))\]

OpenStudy (anonymous):

arccosine and arctan given you the angle given that the result are those numbers

OpenStudy (anonymous):

sin(30 degrees) = 1/2 arcsin(1/2) = 30 degrees

OpenStudy (anonymous):

The first thing to do would be to find the angles from the arc-trig functions

OpenStudy (anonymous):

Can you do those?

OpenStudy (anonymous):

You could also draw two triangles

OpenStudy (anonymous):

 (-3/7) is cos of angle a, (9/4) is the tangent of angle b

OpenStudy (anonymous):

sin(a+b) = sin(a)+cos(b)+cos(a)+sin(b)

OpenStudy (anonymous):

 fixed, the 7 should have been positve